Use a quantifier to make the following true or false, as indicated, where x is a natural number. x - 10 = 10 - x (make false)
A) For x = 10, x - 10 = 10 - x.
B) At least one natural number x exists such that x - 10 = 10 - x.
C) There exists a natural number x such that x - 10 = 10 - x.
D) There is no natural number x such that x - 10 = 10 - x.

Answer by stanbon(75887)

Use a quantifier to make the following true or false, as indicated, where x is a natural number. x - 10 = 10 - x (make false)
A) For x = 10, x - 10 = 10 - x.
True: 0 = 0
-------------------------
B) At least one natural number x exists such that x - 10 = 10 - x.
2x = 20
x = 10
Ans: True
-------------------------
C) There exists a natural number x such that x - 10 = 10 - x.
Ans: True, x = 10
-------------------------
D) There is no natural number x such that x - 10 = 10 - x.
Ans: False
===============================
Cheers,
Stan H.